OpenStudy (anonymous):

Can Someone Explain This To Me? If 2/3p+6=7/6p, what is the value of p

OpenStudy (irishboy123):

this?? \(\frac{2}{3}p+6=\frac{7}{6}p\)

OpenStudy (anonymous):

yes are you using latex I didn't know that worked here

OpenStudy (irishboy123):